Need to find net income. Data are ROE 20%, Debt ratio 55%, Sales 17m, Capital Intensity 1.20 times. I was able to find total debt, total assets, but after that I'm kind of lost...any help?
Answer #1

As capital Intensity is 1.2 and Sales is 17m, we know,

Total Assets = Capital Intensity * Sales
= 17*1.2 = 20.4m.

Now as debt ratio is 55%, ,
value of equity = Total assets (ratio of equity financing)
Value of equity = Total Assets ( 1 - Debt Equity Ratio)
Value of equity = 20.4 ( 1 - 0.55)

= 9.18m.

Now return on equity = (Total Income)/ (Equity value) * 100.

This means,

20% = Total Income / ( 9.18,)

Thus,

Total Net Income = 9.18m * 20%

= 1.836 m (Ans)
Answer #2

SOLUTION :


ROE = Net Income / Equity = 20% = 0.20

Debt ratio = D / A = 55% = 0.55

=> Equity ratio = E / A = 1 - D / A = 1 - 0.55 = 0.45 


Sales = 17m = 17000000

Capital intensity = A / Sales 

=> 1.2 = A / 17000000

=> A = 1.2*17000000 = 20400000 ($) 


=> E = E/A * A = 0.45 * 20400000 = 9180000 ($)


=> Net income = E * ROE = 918000 * 0.20 = 1836000 ($) 


Net income is $1836000 = $1.836 millions (ANSWER)


answered by: Tulsiram Garg